# Realizable Continuous-Space Shields for Safe Reinforcement Learning

### v1

@@ 7 Conclusion
In this paper, we introduce the concept of Proper Shield and develop a novel approach to creating a shield that can handle continuous action spaces while ensuring realizability. Our results demonstrate that this shielding technique effectively guarantees the safety of our reinforcement learning agent in a navigation problem. A significant achievement of our work is a theoretical and empirical integration of liveness requirements into the shielding process. This integration helps mitigate overly conservative behaviors, enabling the agent to recover from loops and repeated actions caused by strict safety specifications, maintaining efficient progress toward its goals. As a limitation, while our definition of a proper shield ensures a response for every state, an exhaustive realizability check may be overly cautious. Some states may not be reachable by a shielded agent and thus not require satisfiability, a factor not considered in our analysis. We plan to investigate this aspect and relax the constraints as a direction for future research.
A major bottleneck in deploying a safe agent for the navigation domain was generating reasonable safety requirements that were also realizable. Looking ahead, we plan to investigate ways to simplify the procedure for generating realizable safety requirements and allow safe RL agents to be more easily deployed across a wide variety of new domains.
Appendix A Algorithm and Implementation Details
@@
Appendix B Training Details
We ran our experiments on a distributed cluster with 255 CPUs and 1T RAM. Each individual training loop was performed on 2 CPUs and 6GB of RAM, for a total wall time of approximately 15 hours. We perform the training with a customized implementation of the Proximal Policy Optimization algorithm (Schulman et al. 2017) from the CleanRL repository [^2] https://github.com/vwxyzjn/cleanrl. Following is the complete list of hyperparameters and a detailed explanation of the state and action spaces.
General Parameters
  • training episodes: 500
  • number of hidden layers: 2
  • size of hidden layers: 32
  • activation function: ReLU
  • parallel environments: 1
  • gamma ($\gamma$): 0.99
  • learning rate: $0.0015$
PPO Parameters
For the advantage estimation and the critic update, we rely on the Generalized Advantage Estimation (GAE) strategy. The update rule follows the guidelines of the OpenAI’s Spinning Up documentation [^3] https://spinningup.openai.com/en/latest/. Following is a complete list of the hyperparameters for our training:
  • memory limit: None
  • update frequency: 4096 steps
  • trajectory reduction strategy: sum
  • actor epochs: 30
  • actor batch numbers: 32
  • critic epochs: 30
  • critic batch numbers: 32
  • critic network size: 2x256
  • PPO clip: 0.2
  • GAE lambda: 0.8
  • target kl-divergence: 0.02
  • max gradient normal: 0.5
  • entropy coefficent: 0.001
  • learning rate annealing: yes
Random Seeds for Reproducibility:
State and Action Spaces:

### v2

@@ 6 Conclusion
In this paper, we introduce the concept of Proper Shield and develop a novel approach to creating a shield that can handle continuous action spaces while ensuring realizability. Our results demonstrate that this shielding technique effectively guarantees the safety of our reinforcement learning agent in a navigation problem. A significant achievement of our work is a theoretical and empirical integration of non-Markovian requirements into the shielding process. This integration helps mitigate overly conservative behaviors, enabling the agent to recover from loops and repeated actions caused by strict safety specifications, maintaining efficient progress toward its goals. Limitations and future directions are discussed in the Appendix G.
Appendix A Different Shielding Techniques
Validate Action Return Safe Action Proper Shield (Def.3.1) Optimization Criteria Continuous State Space Non- Markovian Real Time Naïve Conditions Solver (e.g., Z3) Alshiekh et al. (2018) Rodríguez and Sánchez (2023) This Work
Appendix B Training Details
[…]
Reward Function
For the Mapless Navigation environment, we designed a reward function that incentivize the agent to reach its target position by avoiding the obstacles and additionally provides a penalty to encourage the robot to reach the target with the minimum possible number of steps; more formally:
General Parameters
  • training episodes: 500
  • number of hidden layers: 2
  • size of hidden layers: 32
  • activation function: ReLU
  • parallel environments: 1
  • gamma ($\gamma$): 0.99
  • learning rate: $0.0015$
PPO Parameters
[…]
  • memory limit: None
  • update frequency: 4096 steps
  • trajectory reduction strategy: sum
  • actor epochs: 30
  • actor batch numbers: 32
  • critic epochs: 30
  • critic batch numbers: 32
  • critic network size: 2x256
  • PPO clip: 0.2
  • GAE lambda: 0.8
  • target kl-divergence: 0.02
  • max gradient normal: 0.5
  • entropy coefficent: 0.001
  • learning rate annealing: yes
Random Seeds for Reproducibility:
State and Action Spaces:
Appendix C Algorithm and Implementation Details
@@
Appendix D Anticipation Fragment of LTLt
This appendix extends Subsec. 3.1. In computational logic, a problem is considered decidable if there is an algorithm that can determine the truth of any statement in a finite amount of time. However, reactive synthesis for general Linear Temporal Logic modulo theories (LTLt) is not always decidable. For example, synthesis for expressions like $\square(y>\lhd x)$ is undecidable (Geatti et al., 2023).
[…]
Proof D.1.
Let $\varphi$ be an LTLt formula. Now, let us assume a subformula $\varphi^{\prime}\in\varphi$, where $\varphi^{\prime}:\square[v=t(\lhd v_{1},\lhd v_{2}...)]$, where $v$ is an environment variable, $v_{1},v_{2}...$ are environment or system variables and $t$ is a computable function whose input is $v_{1},v_{2}...$ (e.g., $v_{1}+v_{2}+...$). Also, assume that $v$ appears in predicates $P(v)$ that do not involve other variables (e.g., $v<10$ is valid, but $v<v^{\prime}$ is not valid). Given these conditions, it is easy to see that we can rewrite any subformula in $\varphi$ containing $P(v)$; into a subformula which only depends on the valuations of the past. In particular, we can construct such formula either (1) using the past LTL operator $Y$ (yesterday) Lichtenstein et al. (1985) to construct $\varphi_{NCS}:(P(v_{1},v_{2}...))\wedge\varphi^{\prime\prime}$, where $\varphi^{\prime\prime}$ is the rest of $\varphi$; or (2) by using two temporal layers, i.e., $\varphi_{NCS}:P(v_{1},v_{2}...)\rightarrow\ocircle\varphi^{\prime\prime}$. Therefore, by applying this rewriting (and the rest of $\varphi$ remaining unchanged), any formula containing $\lhd v$ can be transformed into an NCS formula that does not cross state boundaries and we will eventually finish rewriting.
In this paper, our non-Markovian property was not to repeat the visited region, for which the assumptions that Theorem 3.5 needs are enough.
However, what it can happen in the future that we need more expressivity? Say, $\square(v=t(\lhd v_{1},\lhd v_{2}...)\wedge v^{\prime}=t(\lhd v_{1}^{\prime},\lhd v_{2}^{\prime}...)\wedge(v<v^{\prime}))$. We hypothetise that this can also be done (e.g., using the second rewriting method). However, further studying the exact conditions for a formula to belong to the anticipation fragment)is out of the scope of this paper. Last, note that although we rewrite a formula into NCS, the environment is still reactive, because it can be the case that other inputs are free and not determined by the past (e.g., consider a sensor that captures the intensity of the wind).
Appendix E First-Order Theories
First-order theories (Bradley and Manna, 2007) are logical frameworks that use first-order logic (FOL), which extends propositional logic by incorporating quantifiers (i.e., $\forall$ and $\exists$) and variables that can refer to elements in a domain (like integers). A first-order theory $\mathcal{T}$ consists of: (1) a first-order vocabulary used to define terms and predicates, (2) an interpretation that provides meaning to these terms by mapping them to elements in a domain, and (3) an automatic reasoning system that decides if formulas in the theory are valid. A theory’s signature $\Sigma$ consists of constant, function, and predicate symbols, which are the building blocks of formulas, called $\Sigma$-formula, together with logical connectives and quantifiers. The symbols in $\Sigma$ initially have no meaning and are given an interpretation later. A $\Sigma$-formula $\psi$ is satisfiable if there exists an interpretation in $\mathcal{T}$ that makes $\psi$ true. A theory $\mathcal{T}$ is decidable if there is an algorithm that can always determine, in a finite amount of time, whether any $\Sigma$-formula $\psi$ is satisfiable.
Arithmetic Theories.
Arithmetic theories are a particular class of first-order theories which deal with numbers and arithmetic operations. Many of these theories are decidable, though each comes with a different decision procedure and complexity level. Typically, these problems are handled using SMT solvers (Barrett et al., 2024). In this work, we focus on continuous domains represented by Nonlinear Real Arithmetic (NRA) Collins (1974), which is the theory of real numbers with operations like addition and multiplication. The signature for NRA is $\Sigma_{\mathbb{R}}=\{0,k,+,-,=,>,\cdot\}$. For instance, the literal $32x^{2}>1/$ 3 is a valid NRA statement.
Appendix F Safe Reinforcement Learning.
Related Work
Safe reinforcement learning aims to develop policies that not only optimize performance but also respect safety constraints, preventing the agent from taking actions that could lead to harmful or undesirable outcomes. In the literature, this safety aspect can be approached from many different perspectives. A first family of approaches attempts to enforce safer behavior as part of the learning process. An example is constrained reinforcement learning, where the MDP is enriched with a cost function to minimize that represents the safety requirements He et al. (2023); Achiam et al. (2017); Liu et al. (2020); Ray et al. (2019). Other algorithms focus on reward shaping approaches Tessler et al. (2018), policy transfer Simão et al. (2022), behavioral monitors Corsi et al. (2022); Srinivasan et al. (2020), or on techniques that can guarantee safe exploration and data collection phases Marvi and Kiumarsi (2021). However, these approaches are not designed to provide formal guarantees about the final policy, especially in the context of deep learning where neural networks are known to be vulnerable to specific input configurations even when trained with state-of-the-art algorithms Szegedy et al. (2013). Other approaches rely on verification methods to provide guarantees about the behavior of the neural network Katz et al. (2019); Wang et al. (2021); Katz et al. (2017); however, they suffer from significant scalability problems and are limited in neural network topology Liu et al. (2021). Moreover, as many recent papers have shown, as the number of requirements grows, it becomes nearly impossible to find models that formally respect the requirements over the entire input space Corsi et al. (2024a); Amir et al. (2023).
Safe RL Experiments
To further evaluate the necessity of our shielding approach, we conducted additional experiments using a constrained deep reinforcement learning algorithm (Ray et al., 2019), specifically designed to enhance safety. As reported in Tab. 5, while this approach aims to improve the safety of learned policies, we observed that it was still unable to completely eliminate collisions. Moreover, the constrained DRL method provided only marginal improvements over the standard PPO policy, which already demonstrated strong performance in our experiments. These findings underscore the limitations of relying solely on policy optimization, even with safety constraints, to address critical safety requirements. Importantly, our shielding mechanism is not a competing approach but a complementary one, as it can be seamlessly applied on top of any policy, to provide an additional layer of safety during deployment.
Success Collision
$\lambda$-PPO A 0.86 $\pm$ 0.02 0.05 $\pm$ 0.03
$\lambda$-PPO B 0.87 $\pm$ 0.04 0.06 $\pm$ 0.04
$\lambda$-PPO C 0.85 $\pm$ 0.04 0.09 $\pm$ 0.06
$\lambda$-PPO D 0.88 $\pm$ 0.03 0.07 $\pm$ 0.06
$\lambda$-PPO E 0.86 $\pm$ 0.06 0.09 $\pm$ 0.05
Table 5: Comparison of success rate and collision rate achieved using the Lagrangian-PPO algorithm, a constrained reinforcement learning variant of PPO introduced by Ray et al. (2019). The reported models represent the best-performing results selected from a pool of $50$ seeds, and averaged over $5$ runs.
Appendix G Limitations and Future Work
While our definition of a proper shield ensures a response for every state, an exhaustive realizability check may be overly cautious. Some states may be unreachable by a shielded agent and therefore do not require satisfiability, a factor not accounted for in our current analysis. Exploring this aspect and relaxing these constraints represents an important direction for future research.
Another key limitation in deploying a safe agent in the navigation domain lies in generating safety requirements that are both reasonable and realizable. Moving forward, we aim to simplify the process of generating realizable safety requirements, enabling safe RL agents to be more easily deployed across a broader range of domains.